OpenStudy (haziqueaqeel2971):

The Sum of three consecutive even integers is 980. Determine the integers.

OpenStudy (misty1212):

HI!!

satellite73 (satellite73):

guess and check, not algebra try 320, 322, 324 if that doesn’t work, adjust

OpenStudy (haziqueaqeel2971):

Actually, I need help in the thing that we need to use Quadratics to find the numbers.

OpenStudy (misty1212):

that doesnt' work @satellite73 \[320+322+324=966\]

OpenStudy (haziqueaqeel2971):

Like using Quadratic Formula, Discriminant, Completing the square.

satellite73 (satellite73):

yeah @misty1212 i know, i was just giving a start since that is too low, up them

OpenStudy (misty1212):

hmm ok i see but no quadratic formula right?

OpenStudy (haziqueaqeel2971):

I actually don't get what your trying to say.

satellite73 (satellite73):

nope if you want to do it the math teacher way, solve \[n+n+2+n+4=980\]

OpenStudy (misty1212):

isnt that just \(3n+6=980\)?

OpenStudy (haziqueaqeel2971):

Satellite73, then what does N represent?

satellite73 (satellite73):

the first number

OpenStudy (haziqueaqeel2971):

okay, I actually don't understand. It's a thinking question.

satellite73 (satellite73):

consecutive even means like \[324,326,328\]

OpenStudy (haziqueaqeel2971):

Yes yes. But how do you do it?

satellite73 (satellite73):

so if the first one is \(n\) then the second is \(n+2\) and the third is \(n+4\)

OpenStudy (misty1212):

you can solve \[3n+6=980\] right @HaziqueAqeel2971 ?

OpenStudy (haziqueaqeel2971):

Please explain me.

OpenStudy (haziqueaqeel2971):

What do you mean by 3n+6? Is it possible using some quadratic things?

OpenStudy (haziqueaqeel2971):

Using the quadratic formula? Using the completing the square?

satellite73 (satellite73):

quadratic means the degree is 2, like \(x^2+6x=18\)

OpenStudy (haziqueaqeel2971):

Lol, sorry. I am really sorry if you don't understand.

satellite73 (satellite73):

this is just a linear equation, solve in two steps

OpenStudy (haziqueaqeel2971):

Okay, tell me how too. :D

satellite73 (satellite73):

subtract 6, divide by 3

OpenStudy (haziqueaqeel2971):

Ohhh. Then?

satellite73 (satellite73):

then you are donej

satellite73 (satellite73):

\[3n+6=980\] subtract 6 from both sides, then divide both sides by 3, that will give you the first number of the three you are looking for

OpenStudy (haziqueaqeel2971):

Why 3n+6?

satellite73 (satellite73):

explanation is above

OpenStudy (haziqueaqeel2971):

Okay, I saw it. But why did you put 2 + 4?

OpenStudy (mathmale):

I agree with satellite73 here: n+n+2+n+4=980 think what these terms represent: n is the first term. n+2 is the first even number following n. n+4 is the second even number following n. n + n+2 + n+4 is the sum of 3 consecutive even numbers, and this sum is equal to 980. Hope this helps make a little clearer what is happening here.